OpenStudy (anonymous):

(a^3+a^2b/5a)(25/3b+3a)

OpenStudy (anonymous):

\[(3 a+25/(3 b)) (a^3+(a b)/5)\]

OpenStudy (anonymous):

if you write it as \[\frac{a^3+a^2b}{5a}*\frac{25}{3b+3a}=\frac{(a^3+a^2b)(25)}{(5a)(3b+3a)}\] again like the last example look to see how you can simple the top and bottom of the fraction and see if anything cancels. For instance in the numerator, \[(a^3+a^2b)=a^2(a+b)\] and you can also simplify (3a+3b) and then see if things cancel etc etc The final answer you should get is 5a/3

OpenStudy (anonymous):

thanks!!! :)